# Time, Travel, and Energy in the Uniform Dispersion Problem

### v1

@@ III Comparing Makespan, Travel, and Energy
Before proving Proposition III.5, we shall prove a more general claim. A centralized algorithm is an algorithm where a central computer sees, simultaneously, what all the robots in the environment see, and issues commands simultaneously to all of them at the beginning of each time step. If the robots have sensing range $V$, such an algorithm can see any location that is at distance $V$ or less from any robot in the swarm, and control the movements of the robots based on this information. By definition, centralized algorithms require no communication, so they can be parametrized by the visibility $V$ of the robots they control, and the number of persistent state memory bits $S$ that they maintain between time steps.
@@
Definition III.7 helps us capture the idea of a decentralized algorithm simulating a centralized one at the cost of some multiplicative delay $\Delta$ in execution time. Any centralized algorithm ALG can be slowed down into a $\Delta$-delay algorithm ALG’ by simply waiting $\Delta-1$ time steps between each movement command that ALG issues. If a decentralized algorithm simulates ALG’, meaning that at any time step $t$, the decentralized algorithm and ALG’ move robots to the exact same positions, we say that it simulates ALG at delay $\Delta$.
@@
Defining $s$ to be located at $(0,0)$, the algorithm has robots keep track of the $(x,y)$ coordinates of themselves and of other robots’ positions, as well as the coordinates of obstacles. It works as follows: whenever a robot enters a new phase, it spends $\mathcal{O}(V\log n)$ time steps broadcasting a “status message” containing its location, and the coordinates of any previously unreported obstacles it sees. The message requires $\mathcal{O}(V\log n)$ bits because the perimeter of the set of visible locations for each robot consists of at most $4V$ locations, and we report $\log n$ bits per coordinate. Any robot that receives a reported robot or obstacle position then broadcasts it to its neighbours (once per position).
By this propagation process, assuming the visibility graph $G_{t}^{V}$ remains connected, every robot learns the positions of all the other robots during the current phase, and any new obstacle they see, upon receiving at most $n-1$ status update messages. It receives this information during every phase. Hence, by taking into account the robots’ visibility range $V$ and remembering also previously-reported obstacle coordinates, it can know what every robot currently sees. This means it knows everything ALG’ sees at any given time step. Since every robot also has the memory capacity required to execute ALG’ ($\mathcal{S}_{R}$), it can simulate its next step according to ALG’. At every time step, each robot maintains an internal counter. When this counter reaches $\Delta$, the robot moves according to ALG’, resets the counter to 00, updates its phase, and begins entire process anew. Robots only keep track of the parity (even or odd) of their phase, to save memory.
@@
Propositions III.4 and III.5 establish that decentralized swarm-robotic algorithms exist minimizing makespan and travel in all environments. These algorithms assume constant sensing range $V$ and communication bandwidth $B$, and either constant memory capacity $S$ (Proposition III.4) or memory capacity that scales with the size of the environment (Proposition III.5). Let us now show that, as long as the robots’ sensing range is finite, no algorithm exists for minimizing energy in all environments, irrespective of memory capacity and communication bandwidth. In fact, we shall prove that even a centralized algorithm does not exist for minimizing energy:
@@
The proof relies on the construction of a set of environments. Given any algorithm for minimizing energy, ALG, we shall prove it is sub-optimal in at least one environment in the set.
The environment $G(k,r)$ is depicted in Figure 1. It connects $10r$ columns of width $1$. Adjacent columns are spaced $2r$ cells apart. The bottom row is connected to all the columns, and has length $20r^{2}$. All columns but the $1$st and $k$th columns have height $30r^{2}$. The $1$st and $k$th columns have height $130r^{2}+$ 1, and are connected to the top row, as seen in the Figure. The source, $s$, is assumed to be at the bottom left.
@@
At time $t=0$, $A_{1}$ can move either up or right. Since any distinct feature of $G(k,V)$ is at least at distance $V+1$ from $A_{1}$, these directions are indistinguishable to $A_{1}$. Hence, we can assume w.l.o.g. that ALG makes $A_{1}$ take a step up (if it steps right, we simply rotate and reflect $G(k,V)$).
@@
    $v_{1}$ cannot be a vertex in the first column or in the top row except the top vertex of column $k$ or one vertex to its left, as should $v_{1}$ not equal these, settling there would block off the path to parts of the top row via the first column, and force other robots to travel these parts through column $k$. This is sub-optimal, and will cause the total energy to increase beyond $E_{total}^{*}$–a contradiction.
@@
    $v_{1}$ cannot be any vertex in the $k$th column other than the top of the $k$th column, as this would require $A_{1}$ to step downwards.
@@ IV Minimizing Energy in Simply Connected Regions
Figure 2: Possible corners. White vertices are locations in $R$, and blue vertices are walls, i.e., vertices in $R^{c}$.
Figure 3: A hall.
@@
Figure 4: A snapshot from an example run of FCDFS. $s$ is the green circle, black arrows are active robots, black “dots” are settled robots, the blue region is $R^{c}$ and the white region is $R$. The primary direction of each robot is indicated by the direction the arrow representing it is facing. Red arrows point to some of the halls.
Figure 5: Snapshots from a simulation of FCDFS. Rather than block active robots, the settled robots form halls to enable the swarm to explore more of the environment.
@@
Pseudo-code for FCDFS is given in Algorithm 1. The code outlines the Compute step of the Look-Compute-Move cycle. Note that since our robots operate in Look-Compute-Move cycles, when the pseudocode says “step,” we mean that this is the step the robot will take during the Move phase, and not a movement that occurs during computation. In Algorithm 1, we denote by $prev(A)$ the position of a robot $A$ at the beginning of the previous time step, and by $next(A)$ its position at the beginning of the next time step. Section -B shows how FCDFS can be implemented using $5$ bits of persistent state memory (see Algorithm 3).
@@
if $A$ can move in its primary direction then
     Step in the primary direction.
else if $A$ can step in secondary direction then
     Step in the secondary direction.
@@ IV-A Analysis
An articulation point (also known as a separation or cut vertex) is a vertex of a graph whose deletion increases the number of connected components of the graph (i.e. disconnects the graph) [42].
@@
Let $h$ be a hall of a simply connected region $R$. Suppose for contradiction that $h$ is not an articulation point, and let $u$ and $u^{\prime}$ be the neighbours of $h$. Then there is a path from $u$ to $u^{\prime}$ that does not pass through $h$. Let $P$ be this path, and let $P^{\prime}$ be the path from $u$ to $u^{\prime}$ that goes through $h$.
@@
Lemma IV.6 indicates that $R$ can be decomposed into a tree structure $T(R)$ as follows: first, partition $R$ into the distinct connected components $C_{1},C_{2},\ldots,C_{m}$ that result from the deletion of all halls. Letting the vertices of $T(R)$ each represent one of these components, connect $C_{i}$ and $C_{j}$ by an edge if they are both adjacent to the same hall in $R$. We set $C_{1}$ to be the root of the tree, and to be the connected component containing the door vertex $s$.
By Lemma IV.5, assuming our robots correctly stop only at corners, $R(t)$ can in the same manner be decomposed into a tree $T(R(t))$ whose vertices represent connected components $C_{1}(t),C_{2}(t),\ldots$. These components are each a sub-region of some connected component $C_{i}$ of $R$.
@@
We now assume that $v\neq s$. We separate the proof into two cases:
@@
Propositions IV.7 and IV.10 imply that robots take a shortest path in $R$ to their destination. That means that as long as the destination of a robot is not $s$ itself, robots will step away from $s$ one unit of time after they arrive. Until then, this means that robots arrive at $s$ at rate one per two time steps.
@@
IV-B Alternate Optimal Strategies
Figure 7: Robots navigate according to “left hand on wall” until a corner is found.
The key idea of FCDFS to maintain a “geometric invariant”: the environment must remain simply connected after a robot settles. As long as this invariant is maintained, the underlying details of the algorithm can vary. For example, a possible variant of FCDFS that is similarly time- and energy-optimal is shown in Figure 7 (analysis omitted). Rather than stick to their secondary and primary directions, robots move using a “left hand on wall” rule until they hit a corner or a wall.
@@ V Asynchronous FCDFS
These modifications, leaving everything else in our model the same (Section II), result in an asynchronous model of swarms with local signalling. We wish to find a version of FCDFS that works in this asynchronous setting, and to study its performance. FCDFS as implemented in Algorithm 1 relies crucially on the synchronous time scheme to identify halls. This is primarily because, assuming 00 bits of communication bandwidth, robots in our model cannot tell the difference between obstacles (or settled robots) and active robots. In an asynchronous time scheme, the inability to recognize active robots can also cause robots to split away from the trail of robots that tends forms under FCDFS (see Figure 4). Both these issues can be fixed by enabling the robot executing FCDFS to detect active robots. This can be done by giving robots a single bit of communication bandwidth ($B=1$) and having them broadcast ‘$1$’ as long as they are active, indicating that they are active robots (in fact, this requires “less” than a bit of memory, since the robots never need to broadcast ‘00’). Note that, strictly speaking, such robots are no longer ant-like, but they still have very low capability requirements.
An asynchronous version of FCDFS incorporating the above idea is outlined in Algorithm 2. We call this version “AsynchFCDFS” and it is a $(2,1,5)$-algorithm. [^4] The $5$-bit implementation is similar to the synchronous case (Section -B). If we were to optimize, AsynchFCDFS can probably be implemented with even less memory, by taking further advantage of the added bit of communication–but we did not test such implementations.
@@
         Move in the primary direction.
@@
Essentially, Algorithm 2 causes every robot to follow the same path it would under regular FCDFS, by requiring that it wait in place if an active robot is occupying the direction it wants to step in, until that active robot frees up space. This immediately tells us that Algorithm 2 is $T_{total}$ and $T_{max}$-optimal. Our main goal is thus to analyze the energy and makespan cost of the “pauses”. Since our robots’ activation times are now drawn from a random distribution, we will bound these costs probabilistically.

### v2

@@ III Comparing Makespan, Travel, and Energy
Before proving Proposition III.5, we shall prove a more general claim. A centralized algorithm is an algorithm where a central computer sees, simultaneously, what all the robots in the environment see, and issues commands simultaneously to all of them at the beginning of each time step. If the robots have sensing range $V$, such an algorithm can see any location that is at distance $V$ or less from any robot in the swarm, and control the movements of the robots based on this information. By definition, centralized algorithms do not require communication between robots, since a central computer can directly command each robot based on their collective sensory information.
@@
Definition III.7 helps us capture the idea of a decentralized algorithm simulating a centralized one at the cost of some multiplicative delay $\Delta$ in execution time. Intuitively, delay $\Delta$ represents the additional time needed for decentralized robots to gather and propagate information that would be immediately available to a centralized algorithm.
Any centralized algorithm ALG can be slowed down into a $\Delta$-delay algorithm ALG’ by simply waiting $\Delta-1$ time steps between each movement command that ALG issues. If a decentralized algorithm simulates ALG’, meaning that at any time step $t$, the decentralized algorithm and ALG’ move robots to the exact same positions, we say that it simulates ALG at delay $\Delta$.
@@
Defining $s$ to be located at $(0,0)$, the algorithm has robots keep track of the $(x,y)$ coordinates of themselves and of other robots’ positions, as well as the coordinates of obstacles. It works as follows: whenever a robot enters a new phase, it spends $\mathcal{O}(V\log n)$ time steps broadcasting a “status message” containing its location, and the coordinates of any previously unreported obstacles it sees. The message requires $\mathcal{O}(V\log n)$ bits because the perimeter of the set of visible locations for each robot consists of at most $4V$ locations, and when a robot moves, it only needs to report new sightings in the perimeter, since any other location the robot sees at the current time step was visible to it at a previous time step and would have already been reported. Hence, when the robot moves it needs to report the $(x,y)$ coordinates of at most $4V$ obstacles, and each such coordinate can be reported using $\mathcal{O}(\log n)$ bits (as the region consists of $n$ locations thus no coordinate can exceed $n$). Any robot that receives a reported robot or obstacle position then broadcasts it to its neighbours (once per position).
By this propagation process, assuming the visibility graph $G_{t}^{V}$ remains connected, every robot learns the positions of all the other robots during the current phase, and any new obstacle they see, upon receiving at most $n-1$ status update messages. It receives this information during every phase. Hence, by taking into account the robots’ visibility range $V$ and remembering also previously-reported obstacle coordinates, it can know what every robot currently sees. This means it knows everything ALG’ sees at any given time step. Since every robot also has the memory capacity required to execute ALG’ ($\mathcal{S}_{R}$), it can simulate its next step according to ALG’. At every time step, each robot maintains an internal counter up to $\Delta$. When this counter reaches $\Delta$, the robot moves according to ALG’ and resets the counter to 00. This counter is there to get our robots to wait a sufficient number of time steps so that they receive all messages sent out during the current phase. Assuming the phase of the algorithm was $i$, a robot resetting the counter to 00 (note: all robots do this simultaneously) indicates that the algorithm has moved to phase $i+1$. While robots need to count to $\Delta$ within each phase, requiring $\log\Delta$ bits, they need not remember the current phase number, and we save memory by not storing the phase number.
@@
Propositions III.4 and III.5 establish that decentralized swarm-robotic algorithms exist minimizing makespan and travel in all environments. These algorithms assume constant sensing range $V$ and communication bandwidth $B$, and either constant memory capacity $S$ (Proposition III.4) or memory capacity that scales with the size of the environment (Proposition III.5). Let us now show that, as long as the robots’ sensing range is finite, no algorithm exists for minimizing energy in all environments, irrespective of memory capacity and communication bandwidth. In fact, we prove that even a centralized algorithm does not exist for minimizing energy:
@@
The proof relies on the construction of a set of environments. Given any algorithm for minimizing energy, ALG, we prove it is sub-optimal in at least one environment in the set.
The environment $G(k,r)$, parametrized by the integers $r$ (a scaling factor) and $k\in[1,10r]$ (denoting a specific column), is depicted in Figure 1. It contains $10r$ columns of width $1$ spaced $2r$ cells apart, connected by a bottom row of length $20r^{2}$. All columns but the $1$st and $k$th columns have height $30r^{2}$. The $1$st and $k$th columns have height $130r^{2}+$ 1, and are the only columns connected to the top row (see the Figure). The source, $s$, is assumed to be at the bottom left corner.
@@
At time $t=0$, $A_{1}$ is located at the bottom left (since this is where $s$ is), and can move either up or right. Since $A$’s visibility range is $V$, $A$ cannot see any column besides the leftmost column at time $t=0$. This means that $A$ cannot distinguish between $G(k,V)$ and a rotation-and-reflection of $G(k,V)$. Hence, we can assume w.l.o.g. that ALG makes $A_{1}$ take a step up (if it steps right, we simply rotate and reflect $G(k,V)$).
@@
    $v_{1}$ cannot be a vertex in the first column or in the top row except the top vertex of column $k$ or one vertex to its left. Should $v_{1}$ not equal one of these two vertices, there will be vertices in the top row that are to the right of $v_{1}$, but not part of column $k$. Settling at $v_{1}$ would make it impossible for other robots to reach these vertices by going up the first column and then traveling rightward, forcing these robots instead to reach these vertices by going right until the $k$th column, going to the top of column $k$ and then moving left. This “U-turn” is longer than the path that goes up the first column and then goes right, so this will cause the total energy to increase beyond $E_{total}^{*}$–a contradiction.
@@
    $v_{1}$ cannot be any vertex in the $k$th column other than the top of the $k$th column, as this would require $A_{1}$ to step downwards. Stepping downwards results in $A$ taking a sub-optimal “U-turn” path to reach its destination–a contradiction.
@@ IV Minimizing Energy in Simply Connected Regions
Figure 2: Possible corners, denoted by the green circles. White vertices are locations in $R$, and blue vertices are walls, i.e., vertices in $R^{c}$.
Figure 3: A hall, denoted by the green circle..
@@
Figure 4: A snapshot from an example run of FCDFS. $s$ is the green circle, black arrows are active robots, black “dots” are settled robots, the blue region is $R^{c}$ and the white region is $R$. The primary direction of each robot is indicated by the direction the arrow representing it is facing (note this is a state in the algorithm; robots do not have a “heading”). Red arrows point to some of the halls.
Figure 5: Four snapshots from a simulation of FCDFS. Arrows denote primary direction. Rather than block active robots, the settled robots form halls to enable the swarm to explore more of the environment.
@@
Pseudo-code for FCDFS is given in Algorithm 1. The code outlines the Compute step of the Look-Compute-Move cycle. Note that since our robots operate in Look-Compute-Move cycles, when the pseudocode says “step,” we mean that this is the step the robot will take during the Move phase, and not a movement that occurs during computation. In Algorithm 1, we denote by $prev(A)$ the position of robot $A$ at the beginning of the previous time step, and by $prev(prev(A))$ its position two time steps ago. We also denote by $next(A)$ its position at the beginning of the next time step. FCDFS can be implemented with just $5$ bits of persistent memory, but for the sake of a simpler presentation, the pseudocode in Algorithm 1 does not go into the memory management aspect of FCDFS. The complete $5$-bit implementation of FCDFS, including memory management, is deferred to Section IV-B (see Algorithm 2).
@@
if The closest grid cell in $A$’s primary direction is empty then
     Move there.
else if The closest grid cell in $A$’s secondary direction is empty then
     Move there.
@@ IV-A Analysis
An articulation point (also known as a separation or cut vertex) is a vertex of a graph whose deletion increases the number of connected components of the graph (i.e., disconnects the graph) [48].
@@
Let $h$ be a hall of a simply connected region $R$. Suppose for contradiction that $h$ is not an articulation point, and let $u$ and $u^{\prime}$ be the neighbors of $h$. Then there is a path from $u$ to $u^{\prime}$ that does not pass through $h$. Let $P$ be this path, and let $P^{\prime}$ be the path from $u$ to $u^{\prime}$ that goes through $h$.
@@
Lemma IV.6 indicates that $R$ can be decomposed into a tree structure $T(R)$ as follows: first, partition $R$ into the distinct connected components $C_{1},C_{2},\ldots,C_{m}$ that result from the deletion of all halls. Letting the vertices of $T(R)$ each represent one of these components (i.e., each vertex of $T(R)$ represents an entire connected component $C_{i}$ containing possibly many vertices of $R$, not individual vertices of $R$), connect $C_{i}$ and $C_{j}$ by an edge if they are both adjacent to the same hall in $R$. We set $C_{1}$ to be the root of the tree, and to be the connected component containing the door vertex $s$.
By Lemma IV.5, assuming our robots correctly stop only at corners, $R(t)$ can in the same manner be decomposed into a tree $T(R(t))$ whose vertices represent connected components $C_{1}(t),C_{2}(t),\ldots$. These components each represent a sub-region of some connected component $C_{i}$ of $R$. Such tree decompositions will be important for our analysis.
@@
We now assume $v\neq s$, separating the proof into two cases:
@@
Propositions IV.7 and IV.10 imply that robots take a shortest path in $R$ to their destination. That means that as long as the destination of a robot is not $s$ itself, robots will step away from $s$ one unit of time after they arrive. Until then, this means that robots arrive at $s$ at rate one per two time steps (see Figure 7).
Figure 7: Illustration of robot arrival sequence at source $s$. At $t=1$, $A_{1}$ arrives at $s$. At $t=2$, $A_{1}$ moves away from $s$, freeing it up. At $t=3$, $A_{2}$ arrives at $s$ while $A_{1}$ continues forward. This pattern continues, resulting in robots arriving at $s$ at a rate of one per two time steps.
@@
IV-B FCDFS Using 5 Bits of Memory
Algorithm 2 5-bit FCDFS
Let $v$ be the current location of $A$.
if all neighboring vertices of $v$ are occupied then
     Settle.
else if $b_{4}b_{5}=00$ then
     $b_{4}b_{5}\leftarrow 10$
end if
if $A$ cannot move in primary or secondary directions then
     if $v$ has just one unoccupied neighbour then
         Settle.
     else if $(b_{5}=1\land b_{3}+b_{4}=1)$lor $diag(v)$ is unoccupied then
         Settle.
     else
         Set primary direction to obstacle-less direction not equal to $180$-degree rotation of previous direction stepped in (i.e., the neighbour of $v$ we haven’t visited yet; this can be inferred from $b_{1}b_{2}$ and $b_{3}$).
         $b_{4}b_{5}\leftarrow 10$
     end if
end if
if $b_{4}b_{5}$ was not updated at this time step $\triangleright$ i.e., $b_{5}=1$ or time to update $b_{5}$
     $b_{4}b_{5}\leftarrow b_{3}1$
end if
if $A$ can move in its primary direction then
     Move to the closest vertex in the primary direction.
     $b_{3}\leftarrow 0$
else if $A$ can move in secondary direction then
     Move to the closest vertex in the secondary direction.
     $b_{3}\leftarrow 1$
else
     Settle.
end if
As previously discussed, FCDFS can be implemented using 5 bit of memory: see Algorithm 2. In this implementation, a robot’s state is described by bits $b_{1}b_{2}b_{3}b_{4}b_{5}$. All bits are initially 00. $b_{1}b_{2}$ describe the primary direction (one of four), and $b_{3}$ tells us whether the previous step was taken in the primary direction (if $b_{3}=0$) or in the secondary direction (if $b_{3}=1$). $b_{4}b_{5}$ is a counter that is reset to $10$ upon entering a hall or one step after initialization, and thereafter is equal to $*1$, where * is a bit that tells us whether we walked in the primary or secondary direction two steps ago (by copying $b_{3}$). A robot that detects an obstacle at its diagonal interprets its position as a fake hall (i.e., a corner) as long as $b_{5}=1$ and $b_{3}+b_{4}=1$, that is, as long as at least one step passed since the last hall, and our position two steps ago was diagonal to us.
IV-C Alternate Optimal Strategies
Figure 8: Two alternate optimal dispersion strategies. (a) Robots navigate according to “left-hand-on-wall” until a corner is found. (b) Robots stop early and are initialized with random orientations.
The key idea of FCDFS to maintain a “geometric invariant”: the environment must remain simply connected after a robot settles. As long as this invariant is maintained, the underlying details of the algorithm can vary. We depict here two possible variants of FCDFS that are similarly time- and energy-optimal (analysis omitted). In Figure 8(a), rather than stick to their secondary and primary directions, robots move using a “left hand on wall” rule until they hit a corner or a wall. In Figure 8(a), robots settle as soon as they reach a corner (whereas in standard FCDFS they keep moving if possible), and are initialized with random initial orientation. Unlike FCDFS, this variant does not require a compass/dead-reckoning.
@@ V Asynchronous FCDFS
These modifications, leaving everything else in our model the same (Section II), result in an asynchronous model of swarms with local signalling. We wish to find a version of FCDFS that works in this asynchronous setting, and to study its performance. FCDFS as implemented in Algorithm 1 relies crucially on the synchronous time scheme to identify halls. This is primarily because, assuming 00 bits of communication bandwidth, robots in our model cannot tell the difference between obstacles (or settled robots) and active robots. In an asynchronous time scheme, the inability to recognize active robots can also cause robots to split away from the trail of robots that tends to forms under FCDFS (see Figure 4). Both these issues can be fixed by enabling the robot executing FCDFS to detect active robots. This can be done by giving robots a single bit of communication bandwidth ($B=1$) and having them broadcast ‘$1$’ as long as they are active, indicating that they are active robots (in fact, this requires “less” than a bit of memory, since the robots never need to broadcast ‘00’). Note that, strictly speaking, such robots are no longer ant-like, but they still have very low capability requirements.
“AsynchFCDFS” – an asynchronous version of FCDFS incorporating the above idea – is outlined in Algorithm 3. It can be implemented as a $(2,1,5)$-algorithm, with the implementation similar to Section IV-B).
@@
         Move to the closest vertex in the primary direction.
@@
AsynchFCDFS causes robots to arrive at the same locations in the same order they would under synchronous FCDFS, by requiring that robots wait in place if an active robot is occupying the direction they want to step in, until it leaves. This means we retain FCDFS’ no-collision guarantee (Corollary IV.8) despite asynchronicity. It also tells us that Algorithm 3 is $T_{total}$ and $T_{max}$-optimal. Our main goal is thus to analyze the energy and makespan cost of waiting. Since our robots’ activation times are now drawn from a random distribution, we will bound these costs probabilistically.